### v1

# LUMA: A Benchmark Dataset for Learning from Uncertain and Multimodal Data

Abstract
Multimodal Deep Learning enhances decision-making by integrating diverse information sources, such as texts, images, audio, and videos. To develop trustworthy multimodal approaches, it is essential to understand how uncertainty impacts these models. We introduce LUMA, a unique benchmark dataset, featuring audio, image, and textual data from 50 classes, for learning from uncertain and multimodal data. It extends the well-known CIFAR 10/100 dataset with audio samples extracted from three audio corpora, and text data generated using the Gemma-7B Large Language Model (LLM). The LUMA dataset enables the controlled injection of varying types and degrees of uncertainty to achieve and tailor specific experiments and benchmarking initiatives. LUMA is also available as a Python package including the functions for generating multiple variants of the dataset with controlling the diversity of the data, the amount of noise for each modality, and adding out-of-distribution samples. A baseline pre-trained model is also provided alongside three uncertainty quantification methods: Monte-Carlo Dropout, Deep Ensemble, and Reliable Conflictive Multi-View Learning. This comprehensive dataset and its tools are intended to promote and support the development and benchmarking of trustworthy and robust multimodal deep learning approaches.
1 Introduction
In recent years, the use of Machine Learning and Deep Learning has surged across various fields, driving advancements in data analysis and decision-making. In domains such as healthcare, autonomous driving, and finance, information is distributed across multiple modalities including audio, video, text, and images. To better understand the data and improve decision-making capabilities, it is crucial for deep learning models to integrate diverse, multimodal sources of information. Multimodal Deep Learning (MDL) addresses this need and improves the capabilities of uni-modal networks [4, 20, 36, 22].
Another important consideration for deploying deep learning models in safety critical fields is trustworthiness. Traditional deep learning models are often overconfident in their predictions [1], which can lead to catastrophic results in areas such as healthcare or autonomous driving. Although various techniques for uncertainty quantification have been proposed to measure the level of uncertainty in data and model, this remains an open and challenging area. More research and robust benchmarks are needed to advance the field of uncertainty quantification in deep learning [18, 27].
In probabilistic modeling, uncertainty is usually divided into aleatoric (data) and epistemic (model) uncertainties [15]. Aleatoric uncertainty refers to the uncertainty in the data due to inherent noise. It is impossible to reduce the amount of aleatoric uncertainty with additional data (hence, it is also often called irreducible uncertainty). Epistemic uncertainty is the uncertainty in model parameters, due to lack of data, hence, it can be reduced with additional data samples. Epistemic uncertainty is also usually high for Out-of-distribution (OOD) data, and is commonly used for OOD detection.
Multimodal uncertainty quantification (MUQ) is a relatively new research area that adapts uncertainty quantification approaches to multimodal deep learning problems, aiming to enhance the trustworthiness of these models [12]. Due to the unsupervised nature of uncertainty quantification, where the exact extent of uncertainty in the data and the model is unknown, analyzing and benchmarking proposed UQ methods is challenging. Current multimodal datasets used for benchmarking state-of-the-art models in multimodal uncertainty quantification [9, 12, 13, 23, 37] lack the ability to inject a controlled amount and various types of uncertainties for each modality. This limitation hinders the comprehensive benchmarking of MUQ techniques, which is essential for developing trustworthy and robust multimodal deep learning approaches.
To address this issue, we present LUMA (Learning from Uncertain and Multimodal dAta), a multimodal dataset designed for benchmarking multimodal learning from uncertain data. This dataset currently consists of 101,000 images, 135,096 audio recordings, and 62,875 text passages, representing a total of around 3 GB, with a Python toolkit for uncertainty generation. The modalities are carefully aligned, ensuring that each text passage is related to the object in the corresponding image, and each audio recording is the pronunciation of the object label in the image. The toolkit allows the injection of aleatoric and epistemic uncertainties in a controlled and parameterized way into each modality specifically.
To summarize, our contributions are as follows:
  1. 1.
    We propose LUMA [^1] https://huggingface.co/datasets/bezirganyan/LUMA; , a benchmarking dataset for learning from uncertain and multimodal data. It includes audio, image, and textual modalities across 50 distinct classes. We compiled the images from the CIFAR 10/100 dataset [19], extracted, validated, and associated the corresponding audio samples from three diverse audio corpora, and generated the related text modality utilizing Gemma-7B Instruct [26] Large Language Model (LLM). We also performed additional bias analysis of the dataset. Each generated version of the dataset consists of 600 data records per class (500 for training, and 100 for testing) belonging to 42 classes, and 3,859 OOD data points, belonging to the remaining 8 classes.
  2. 2.
    We offer a Python package [^2] https://github.com/bezirganyan/LUMA that generates dataset samples with varying levels of noise and uncertainty. The uncertainty generator can effectively increase aleatoric uncertainty in the data and epistemic uncertainty in the model.
  3. 3.
    Finally, we provide baseline models including three different uncertainly quantification methods (Monte-Carlo Dropout [7], Deep Ensemble [21], Reliable Conflictive Multi-View Learning [37]), to serve as a starting point for benchmarking.
2 Limitations of current Datasets for MDL benchmarking
In practice, we often don’t know the extent of inherent uncertainties in the data or how accurately they represent the real-world data space. This often makes it hard to evaluate how well uncertainty quantification algorithms work. Moreover, deep learning algorithms may behave differently under different amount of uncertainties (i.e., the robustness to noise may vary). Thus, it may be beneficial to inject additional amount of noise in the data, and observe the change in uncertainty metrics and the performance of the models. Since approaches to quantify different types of uncertainty vary, it is beneficial to have options for injecting various types of uncertainties.
Several datasets are used in multimodal uncertainty quantification settings. A notable line of work [9, 12, 13] has employed datasets such as HandWritten [^3] https://archive.ics.uci.edu/ml/datasets/Multiple+Features, CUB [^4] http://www.vision.caltech.edu/visipedia/CUB-200.html, Scene15555https://serre-lab.clps.brown.edu/resource/hmdb-a-largehuman-motion-database, and Caltech101  [^6] https://data.caltech.edu/records/mzrjq-6wc02. These datasets typically extract different features from unimodal sources to create a multi-view setup. While they have been instrumental, they primarily repurpose unimodal data for multimodal tasks, underscoring the need for more comprehensive and inherently multimodal datasets to better evaluate uncertainty in deep learning models.
Furthermore, the current approaches that introduce uncertainty in the data [9, 12, 13] add Gaussian noise to the views or the extracted features. While Gaussian noise does increase uncertainty, it does not accurately reflect the noise that can be found in real-world datasets and this process lacks fine-grained control over the type of uncertainty being injected.
Additionally, how different modalities’ uncertainties interact significantly impacts the overall multimodal uncertainty. When both modalities encode redundant information, the total uncertainty might not decrease. Conversely, conflicting information can lead to increased uncertainty, while complementary information can reduce it. A deeper understanding of these phenomena is crucial. Fine-grained control over individual modalities’ uncertainties opens the way for more theoretical research based on empirical observations.
To better understand and analyze uncertain multimodal data, as well as to debug and benchmark uncertainty quantification techniques in the multimodal learning context, we propose a dataset accompanied by an uncertainty generator package. This package includes various techniques for injecting uncertainty, such as controlling data diversity, adding different types of real-world noise, randomly switching labels to their closest class, and injecting out-of-distribution (OOD) data.
3 LUMA Dataset
In this section, we introduce LUMA, a dataset composed of an extensible list of modalities including image, audio, and text modalities, collected from various sources.
3.1 Image modality
For the image modality, our priority was to choose a relatively simple yet well-known dataset, where we could have the option to manually increase the degree of uncertainty. For that purpose, we chose CIFAR-100 and CIFAR-10 [19] datasets since they are well-known datasets of small 32x32 images, with lots of baseline models. 42 classes were chosen so that after aligning with the other modalities, we would have at least 600 samples in each class per modality. The 600 threshold was selected based on the number of images per class in the CIFAR-100 dataset. We took another 8 classes, which had less than 600 samples after aligning with other modalities, as OOD samples. In total, we took 25,200 images as train/test data, and 3,859 images as OOD data (see the image collection pipeline in Figure 1).
Figure 1: Image collection pipeline
Aside from the main dataset, as described in the section 3.4.1, another priority was to understand the behaviors of models under different levels of data diversity. To achieve this, we decided to sample 600 data points with different level of diversity from the bigger set CIFAR-10/100. However, in CIFAR-100 dataset, there are no more than 600 samples per class. We alleviated this issue with including images generated with EDM Diffusion-based generative model [^7] Retrieved generated samples from https://github.com/wzekai99/DM-Improves-AT[14]. We chose EDM-generated images, since the generated samples were already available, and Zheng et al. [38] showed that augmenting CIFAR-10 data with EDM-generated samples significantly improves the classification accuracy.
3.2 Audio modality
For audio modality, the diversity of accent in the pronunciation was an important factor to be considered and we collected samples, where different people would pronounce the corresponding class label. For this task, we used three audio/text parallel corpora, and extracted the desired audio segments. More specifically, we used The Spoken Wikipedia [16], LibriSpeech [28], and Mozilla Common Voice [2] corpora. The audio collection pipeline is depicted in Figure 2.
Figure 2: Audio collection, extraction and validation pipeline
The Spoken Wikipedia is a collection of hundreds hours of phoneme-level aligned audio, where volunteer readers are reading various Wikipedia articles. We used these alignments to extract all the instances of audio segments that pronounced one of the CIFAR-10/100 classes.
The LibriSpeech dataset is a corpus of 1,000 hours of English speech, derived from audiobooks from the LibriVox [^8] https://librivox.org/ project, which is a collection of public domain audiobooks. Unfortunately, LibriSpeech doesn’t provide word-level alignment, hence, we used force-aligned alignments [^9] https://github.com/CorentinJ/librispeech-alignments generated with the Montreal Forced Aligner [25]. Similarly to The Spoken Wikipedia, we looked up the CIFAR-10/100 labels in forced aligned textual data, and extracted the corresponding audio segments.
The Mozilla Common Voice corpora, is a crowdsourced open-source collection of voices by volunteer contributors from around the world. Like LibriSpeech, Mozilla Common Voice also doesn’t provide word-level alignments, hence, we again used forced aligned alignments [^10] https://github.com/JRMeyer/common-voice-forced-alignments, and extracted the relevant audio samples.
73 Additional recordings of pronunciations belonging to 4 classes ("roses", "telephone", "whale", "wolf") were voluntarily contributed by our colleagues, which were anonymized, trimmed, and added to the dataset.
From these corpora, we used the following rule to extract the samples. First, we extended our class label set with a superset that also contains the plural forms of the words (i.e., for the audio track “horse”, the audio track “horses” was added to the set), then we iterated over all aligned transcripts, and for any word included in the formed set, we extracted the corresponding audio sample. We considered the plural forms, since we believe that an extra "s" or "es" does not change the pronunciation of the words much. We did not consider the plural forms if it requires audible changes to the word root (i.e., mouse - mice). The extraction algorithm can be found in the supplementary material.
Since most of the audio data is collected from forced alignments, it is possible to have misaligned audio segments, which could introduce additional noise to the dataset. Moreover, since part of the audio samples are from voluntary contributions, there can be very noisy samples, which are hard to interpret, or audio samples with a strong accent, which again can be hard to interpret. To remove such extreme cases in audio samples, we performed an automatic validation of the samples. Then, we filtered out the false negatives with manual validation for the negative predictions.
The automatic validation was achieved with the OpenAI’s Whisper Large V3 model [31] for audio transcription, and transcribed the extracted audio samples. If the transcription corresponded to the class label (or its plural form), then we considered the sample as valid. Otherwise, the sample was sent for manual validation. Because of the huge output space of the Whisper Large V3 model, the probability of false positives is quite low, so we did not perform a manual validation for positive predictions. To summarize, we validated 130,069 out of 178,123 data samples with automatic validation, and we performed a manual validation for the remaining samples.
For manual validation, we decided to check only the classes, which did not have more than 800 samples (to be able to sample 600 samples with different degrees of diversity, as described in Section 3.4.1). Hence, we filtered 8,372 samples, and scheduled them for manual labeling. We opted for Label Studio [33] to build the labeling interface (See the supplementary material for the annotation interface). The interface provided the audio sample, with the prompt "Is the audio saying the word below? (An extra ’s’ or ’es’ in the pronunciation is okay.)" and answer options of "Yes" or "No". We asked our colleagues (M.Sc. and PhD Students, and Professors) with advanced to fluent English knowledge to annotate the samples.
In total, we collected 2 annotations per sample, from 17 annotators. We got 71.61% of annotation agreement, and accepted 5,027 samples, where both annotators confirmed the validity of the sample. Hence, we took the 42 classes that had more than 600 validated samples (automatically and manually) as training/test data, and we took the remaining 8 classes as OOD data. In total, the auto-validated and manually validated audio samples combined, LUMA has 135,096 audio samples. The final distribution of audio data across classes can be seen in the supplementary material.
3.3 Text Modality
For text modality, the main constraint was that the text segments had to talk about the subject of the images. For that, we decided to employ a generative model, and generate text segments about the class label. We utilized Google’s Gemma-7B Instruct model [26] to generate more than 1,200 texts samples per class, using 13 different prompts. Here is an example of a prompts used for generation (please find the full list of the prompts in the supplementary material).
"You are explaining a five year old child what the word <word> means. Use very simple and explanatory language, so the kid will understand the meaning of the word <word>. Tone: Casual, complexity: simple"
Gemma-7B Instruct was chosen, since according to their technical report [26], it outperforms other open LLMs with similar size, in 11 out of 18 tasks. Moreover, in our experiments, it provided better answers to our prompts compared with Mistral-7B [11].
Figure 3: Text generation and validation pipeline
To validate that the generated texts accurately represent the labels, we masked all label occurrences in the text and fed the masked text back into the Gemma-7B Instruct model, asking it to classify the text into one of the labels (see the prompt in the supplementary material). Based on the prediction of the model, if the prediction matched the ground truth label, we accepted the sample as validated. In total, we accepted 55,953 text samples.
After manually analyzing some of the generated texts, we noticed that there were samples with offensive biases and stereotypes (Some examples are included in the supplementary material). Particularly, we noticed a lots of gender bias for classes "man", "woman", "boy" and "girl". To find the proportion of the biased data, we asked the Gemma model to find out if the given text contains gender, racial, religious, or cultural biases. We found out, that indeed, the aforementioned 4 classes have a huge amount of gender bias (see the bias detection statistics in the supplementary material). Our hypothesis is, that describing a man or woman in an unbiased way is a challenging task for LLM models (as well as for humans), which are trained on unbalanced data [17]. For that purpose, we reconstruct the prompts (see in the supplementary material), to explicitly provide topics and keywords with occupations, which will minimize the bias. We then rerun the bias detection prompt, which found fewer biased samples and allowed us to filter enough samples to be included in the final dataset.
Since textual data was generated using an LLM, we recognize that the dataset may contain factual inaccuracies, or biases, but our aim is to offer a benchmark to study uncertainty quantification in multimodal classification settings. LUMA shall not be used as a source of knowledge or information.
3.4 Dataset compilation
Based on the collected samples from the 3 modalities (image, text, audio), we wanted to compile a dataset with little uncertainties, and later, provide tool to inject uncertainties on demand. Our priority was to propose several options for uncertainty control and parameter setting: data diversity, sample noise, label noise, OOD injection.
3.4.1 Data Diversity
Figure 4: t-SNE [35] visualization of audio data points from class "man" (in gray), and sampled points with different diversity parameter $k$. The higher is the value of $k$, the more concentrated (less diverse) the points are.
With a fixed number of data points, increasing the diversity of the data enhances the information passed to the model, thereby it shall reduce the epistemic uncertainty. Conversely, when samples are concentrated at a single point in the latent space, they encode less information, which shall lead to greater epistemic uncertainty in areas where data is scarce (Figure 5). Hence, controlling the diversity of the data allows us to study the behavior of epistemic uncertainties under varying amounts of information.
Figure 5: t-SNE [35] visualization of audio data points for all classes, sampled with different diversity parameter $k$. With higher $k$ we have more concentrated samples, and more separation between classes. The diversity can similarly be controlled for the other modalities.
To control the diversity, we extract deep features from each modality (Wav2Vec [3] for audio, BERT [6] for text and VGG-11 [32] for images), and compute the inverse distance of each sample to the center (mean vector) of its class, raised to the power of $k$:
$D_{i}=\frac{1}{\left\|F_{i}-\frac{1}{|C|}\sum_{j\in C}F_{j}\right\|_{2}^{k}},\ \ i\in C,$ (1)
where $F$ represents the deep feature vectors extracted from the samples, $C$ is the set of data indices belonging to the class, and ||$\cdot$|| ⋅ | measures the cardinality of the set. Then, having the inverse distances, we sample points from categorical distribution $x_{n}\sim Categorical(D)$. In Eq. 1, $k$ is the variable controlling the diversity. If $k=0$ the sampling is uniform. The bigger $k$, the higher probability of selection will be applied to the samples closer to the center.
Having sufficient samples in image and text modalities, our bottleneck was the number of samples in the audio modality. Since in the 42 in-distribution classes, around 70% have more than 900 audio samples, we considered this enough for diversity control.
Sample Noise: We want to have an option to inject controlled amount of noise in the data. This may reduce the information in each data sample, and increase the classification difficulty. With our hypothesis, this may affect both epistemic and aleatoric uncertainty degrees. This type of noise can also be very beneficial for estimating the model robustness to noise. We apply different types of noise to each modality.
For audio modality, we added background noise from ESC-50 dataset [30] to each sample, using the audiomentations [^11] https://github.com/iver56/audiomentations library. The amount of the minimum and maximum signal-to-noise ratio, as well as the proportion of the noisy data is set as a hyper-parameter.
For text modality, we utilize the nlpaug [24] library, to add different types of noise. The user has option to choose a subset of noise types from: 1) Keyboard noise that simulates keyboard distance error; 2) OCR noise that simulates OCR engine noise; 3) Random character noise to insert, substitute, or delete random characters; 4) Antonym noise to swap random words with their antonyms; 5) Random word noise to insert, substitute, or delete random words; 6) Spelling noise to add spelling mistakes according to spelling mistake dictionary; and 7) Back-translation noise to translate the text to another language, and then translate back to English. The parameters of these noise types can be specified by the user, and are transferred to the nlpaug library for adding the specific type of noise.
For image modality, we added different types of noise suggested and implemented by Hendrycks and Dietterich [10]. 15 perturbations are included such as: adding Gaussian noise, shot noise, impulse noise, defocus blur, frosted glass blur, motion blur, zoom blur, snow, frost, fog, changing the brightness, contrast, elasticity, pixelating, and JPEG compressing.
Label Noise: Another way to insert aleatoric uncertainty is to add label noise in the data (i.e., randomly switch the labels of some samples). Since the uncertainty induced by this type of noise cannot be reduced with additional amount of data, this noise shall increase the aleatoric uncertainty of the model. To insert this type of noise, we choose samples at random (according to a probability defined by the user), then, based on their deep features (Wav2Vec [3] for audio, BERT [6] for text and VGG-11 [32] for images), we find the average distance to the 5 nearest points from each class, and assign the label of the class that has the closest mean distance.
OOD Injection: Ideally, the models shall be uncertain on data points from unknown distribution (i.e., distribution they haven’t been trained on). In the literature, often the OOD samples are taken from another dataset, which can simplify the problem, because such samples are far from the training data. For this matter, we kept a separate set of samples from the same dataset, but belonging to classes that are not present in the training data, as OOD samples.
4 Baseline models with Uncertainty Quantification
We develop a baseline model with three different uncertainty quantification algorithms, to serve as a starting point for other research and benchmarking initiatives. For the sake of simplicity, we choose late or decision fusion approaches, where we have classification networks for each modality, and then fuse their decision by simple averaging the output logits. The implementation details of the architectures can be found in the supplementary material.
For uncertainty quantification, we implemented 3 approaches: Monte Carlo Dropout (MCD) [7], Deep Ensemble (DE) [21], Reliable Conflictive Multi-View Learning (RCML) [37]. In Monte Carlo Dropout and Deep Ensembles, we use the aleatoric entropy and the epistemic entropy as as uncertainties measures $H_{\text{Ale }}(y\mid\mathbf{x})=\operatorname{entropy}\left(p_{\text{Ale }}(y\mid\mathbf{x})\right)$ and $H_{\mathrm{Epi}}(y\mid\mathbf{x})=\operatorname{entropy}\left(p_{Epi}(y\mid\mathbf{x})\right)$, where $p_{Epi}$ and $p_{Ale}$ are the probabilities obtained according to [34]. In RCML, we measure the aleatoric uncertainty with the expected entropy such as:
$\mathbb{E}_{p(\boldsymbol{\pi}\mid\mathbf{x},\hat{\boldsymbol{\theta}})}[H[P(y\mid\boldsymbol{\pi})]]=-\sum_{k=1}^{K}\frac{\alpha_{k}}{\alpha_{0}}\left(\psi\left(\alpha_{k}+1\right)-\psi\left(\alpha_{0}+1\right)\right),$ (2)
where $\alpha_{k}$ is the $k$-th concentration parameter of the Dirichlet distribution, and $\alpha_{0}$ is the sum of all concentration parameters. $\psi$ is the digamma function. As a measure for epistemic uncertainty, we take $\frac{N}{\alpha_{0}}$, where $N$ is the number of classes.
We evaluate the measures of accuracy and uncertainty of the models on the clean dataset, the dataset with reduced diversity ($\searrow$ Diversity), the dataset with increased sample noise ($\searrow$ Sample Noise), and the dataset with switched label $\nearrow$ Label Noise). The results are summarized in Table 1. (For noise generation parameters, please refer to the supplementary material). Since in RCML the uncertainty is quantified differently compared with MCD and DE, we cannot directly compare their values, and relative changes to the two type of uncertainty compared to the clean dataset is reported in the table.
Table 1: Results for UQ with baseline models. The absolute values are reported for clean dataset, and changes in percentages relative to clean dataset are reported for the noisy versions of LUMA dataset.
Method Clean $\searrow$ Diversity $\nearrow$ Label Noise $\nearrow$ Sample Noise
Ale. Epi. Ale. Epi. Ale. Epi. Ale. Epi.
MCD Image 1.00 1.03 -15.73% -11.66% +59.20% +54.51% +4.44% +2.18%
MCD Audio 0.52 0.70 -5.54% +2.16% +96.63% +54.49% +23.12% +14.40%
MCD Text 0.37 1.01 -3.91% -2.62% +93.59% +2.41% +64.96% -2.03%
MCD Multi. 0.26 0.78 -8.52% -1.21% +122.44% +11.60% +59.14% +9.89%
DE Image 1.45 1.40 -37.49% -8.54% -7.43% +0.24% -18.46% -3.22%
DE Audio 0.56 0.99 -27.39% -3.34% +156.40% +50.43% +70.26% +34.41%
DE Text 0.42 1.01 +5.02% -6.15% +81.26% -0.51% +62.24% -7.11%
DE Multi. 0.31 0.82 -22.80% -3.40% +115.15% +20.62% +45.97% +5.54%
RCML Multi. 1.99 0.43 +8.34% +16.16% +64.72% +106.16% +36.19% +58.21%
Table 2: OOD Detection AUC Values for Different Methods
Method MCD DE RCML Multi.
Image Audio Text Multi. Image Audio Text Multi.
AUC 0.54 0.47 0.53 0.50 0.54 0.49 0.54 0.50 0.91
As we can observe in the table, in most cases, adding label and sample noises effectively increases the epistemic and aleatoric uncertainties. Interestingly, in most MCD and DE models, the uncertainty decreases when they are trained on data with lower diversity. This may indicate that these approaches fail to recognize data points outside their training distribution, which we will further investigate with the OOD detection task.
We evaluate AUC score for OOD detection by the networks based on the epistemic uncertainty. The results are summarized in Table 2. We can see that Monte Carlo Dropout and Deep Ensembles fail to provide epistemic uncertainty values suitable for OOD detection in LUMA dataset, with a poor performance of approximately 0.5 AUC value. On the other hand, the RCML achieves an outstanding AUC score of 0.91, indicating that the epistemic uncertainty values quantified with this method can be effectively used for OOD detection.
In conclusion, the performance of Monte Carlo Dropout and Deep Ensembles indicates a limitation in their suitability for OOD detection in LUMA dataset. This suggests new avenues for further exploratory research to leverage uncertainty estimation for robust detection of out-of-distribution samples. Furthermore, the observed disparities highlight the necessity for a comprehensive benchmarking effort on LUMA dataset, encompassing a broader array of state-of-the-art methodologies.
5 Conclusion
In this paper, we propose LUMA, a new multimodal dataset for learning from uncertain and multimodal data and benchmarking. It includes image, audio and text modalities and a Python package for compiling different versions of the dataset with various amounts and types of uncertainty and noise.
While LUMA multimodal dataset and tools offer resources for effectively studying and extensively benchmarking multimodal uncertainty quantification, it’s important to acknowledge some limitations. 1) Potential bias in the text modality: The text data was generated by a large language model, which can introduce biases and errors. Despite manual and automatic efforts to mitigate such bias, the texts may still contain undetected biases or errors. Therefore, we emphasize that this dataset is strictly for research on multimodal learning and uncertainty quantification. 2) Limited audio samples: While all 42 in-distribution classes have at least 600 audio samples, our diversity reduction algorithm might not perform optimally for classes with only slightly more than 600 samples. This limitation could hinder studies focusing on the audio modality.
The dataset can be easily extended with additional modalities and augmented with more data samples. The open-source nature of the data compilation pipeline and code for uncertainty and noise generation facilitates the integration of new contributions from the community to promote multimodal uncertainty studies and benchmarking initatives.
Acknowledgments and Disclosure of Funding
We thank our colleagues and friends for contributing to the audio validation process, and to the HumanSignal team for providing an academic license for Label Studio.
This work is funded by the Institut Archimedes doctoral program of Aix-Marseille University (AMU).
Supplementary Material
In this section we present the supplementary material for LUMA dataset.
The supplementary material contains:
  • A - Details about Data Collection
  • B - Details about the included baseline architectures and datasets
  • C - Datasheet for LUMA dataset
Appendix A Details about Data Collection
A.1 Image Modality
A.1.1 Licenses of the source datasets
As mentioned in Section 3.2 of the main paper, we use the CIFAR-100 and CIFAR-10 [19] datasets and EDM [14] generated images from the same distribution for the image modality.
Unfortunately, we failed to find an explicit license for the CIFAR-10/100 datasets, hence, we did not include the images in our dataset repository, and instead, allow the users to download them directly using the provided LUMA dataset compilation and uncertainty generation tool.
The EDM Generated images retrieved from https://huggingface.co/datasets/P2333/DM-Improves-AT are published under the Apache-2.0 license.
A.2 Audio Modality
In this section, we will present some additional information regarding data collection process of audio modality.
A.2.1 Licenses of the source datasets
As mentioned in Section 3.2 of the main paper, we collected audio data from 3 sources: The Spoken Wikipedia [16], LibriSpeech [28], and Mozilla Common Voice [2] corpora.
The Spoken Wikipedia dataset is published under the CC BY-SA 4.0 license. Following the requirements of this license, we also distribute our dataset under the same license. The audio samples extracted from this dataset are located in the https://huggingface.co/datasets/bezirganyan/LUMA/tree/main/audio/sw_audio directory in LUMA dataset.
The LibriSpeech dataset is published under the CC BY 4.0 license. The audio samples extracted from this dataset are located in the https://huggingface.co/datasets/bezirganyan/LUMA/tree/main/audio/ls_audio directory in LUMA dataset.
The Mozilla Common Voice dataset is published under the CC0 (public domain) license. The audio samples extracted from this dataset are located in the https://huggingface.co/datasets/bezirganyan/LUMA/tree/main/audio/cv_audio directory in LUMA dataset.
A.2.2 Audio extraction and validation
To extract the audio pronunciation of each word as described in Section 3.2, we follow the Algorithm 1. First, the algorithm iterates through all the words in the transcripts. If a word matches one of the labels or their plural forms, it is added to the extraction_words list (Lines 3-10), along with its corresponding time steps and audio path. Next, the algorithm processes each word in this list, using the extract_audio function to extract the relevant audio segments. These segments are then added to the extracted_audios list, which is returned as output of the algorithm (Lines 11-17).
Algorithm 1 Audio Extraction Algorithm
1:  Input: labels $\leftarrow$ list of the CIFAR-100/10 labels
2:  Input: transcripts_paths $\leftarrow$ list of (transcript, audio_path) tuples
3:  extraction_words $\leftarrow$ empty list
4:  for transcript, audio_path in transcripts_paths do
5:     for word in transcript do
6:        if word in labels or word + ’s’ in labels or word + ’es’ in labels then
7:           extraction_words.append((word, transcript[word].start_time, transcript[word].end_time, audio_path))
8:        end if
9:     end for
10:  end for
11:  extracted_audios $\leftarrow$ list
12:  for entry in extraction_words do
13:     word, start_time, end_time, path $\leftarrow$ entry
14:     audio_segment $\leftarrow$ extract_audio(path, start_time, end_time)
15:     extracted_audios.append(audio_segment)
16:  end for
17:  return  extracted_audios
Then, we perform automatic validation with the OpenAI’s Whisper Large V3 model [31], and on the subset of the negative predictions of the model we perform a manual validation. For the manual validation, we have collected manual annotations thanks to the efforts of our volunteering colleagues and friends. To build the annotation interface, we opted for the Label Studio [33], an open-source data labeling platform. You can see the screenshot of the annotation interface in Figure 6. In total, we collected 16,744 manual annotations from 17 human annotators over a period of 2 months. Each sample was annotated by 2 annotators, and only annotations with 100% agreement were accepted.
After automatic and manual validation, we have 135,096 audio samples with the class distribution shown in Figure 7.
Figure 6: The labeling interface using Label Studio
Figure 7: Number of validated audio samples per class. We will include the classes with higher than 600 samples as in-distribution data, and others as out-of-distribution data.
A.3 Text Modality
For the text modality, as described in the Section 3.3, we employed Google’s Gemma-7B Instruct model [26] to generate more than 1,200 text samples per class, using 13 different prompts. The 13 prompts are as follows:
"You are talking with your friend about some topic. Use the word <word> in a sentence with your friend. Use casual language. Tone: Casual / Conversational, length: short",
"You are the prime minister of the United Kingdom. During a press conference you are asked a question about <word>. Give a sentence from that press conference mentioning the word <word>. Tone: Formal, length: medium",
"You are explaining a five year old child what the word <word> means. Use very simple and explanatory language, so the kid will understand the meaning of the word <word>. Tone: Casual, complexity: simple",
"Imagine you are writing a science fiction book. Write a conversation from that book mentioning the <word>.",
"You are the editor in a mainstream journal. Write a sentence from a news article about a <word> in your journal that mentions the word <word>.",
"You are a teenager writing a post in Facebook about <word>. Write the post about the experience you had with the <word>.",
"You are playing a word describing game with your friend. The word is <word>, and you shall describe it without mentioning the word itself, so your friend will guess it. Explain it to him clearly in a simple language.",
"Think of something else that shares similar characteristics or functions with the <word>. Draw comparisons or use analogies between that other word and the <word>.",
"Place the word <word> within historical context. How would you describe it in relation to its origins, evolution, or significant historical events? Be creative in your description.",
"Consider how the word ’<word>’ is depicted or referenced in popular culture, literature, or media. Describe it by referencing these cultural elements.",
"Pretend you are a character who sees a <word> for the first time in your life. Describe it from the characters perspective, considering their background, personality, and knowledge.",
"Write a 4 line small poem about the word <word>. Be creative, and use casual tone for the poem.",
"You are a musician composing a song inspired by <word>. Write the lyrics to the song, capturing the mood, emotions, and imagery associated with <word>. Use rhythm and melody to convey the essence of <word> in your music."
The <word> was replaced with the class labels. After generating the text data, to check if the generated text has any bias or stereotype, we again used the Gemma-7B Instruct model with the following prompt:
Your job is to identify biases in texts. You will be given a text, and you need to classify it into one of: [gender_bias, cultural_bias, racial_bias, religous_bias, no_bias].
[…]
The <text> was replaced with the text that needed to be checked. The number of biased and unbiased texts in each class according to Gemma model can be found in Figure 8. As we can see, the labels man, woman, boy, and girl have high amount of biased texts. Some examples of identified biased texts are:
A woman is a grown-up person who has a soft, nurturing personality. She usually takes care of her family and friends, and sometimes works outside the home. Women are strong and smart, they can do many things that men can do.
Man, a force of might,
A guardian, protector, and light.
With strength and wisdom, they stand tall,
Strong and proud, answering the call.
A tool in a toolbox is an efficient and valuable asset that aids in various tasks. Similarly, a man is also a valuable asset to any group or society. Just like a tool in a toolbox, a mans capabilities are tailored to fulfill different roles and functions, making him an essential component of any endeavor.
Figure 8: The amount of texts with different biases according to Gemma-7B Instruct model.
Figure 9: The amount of texts with different biases according to Gemma-7B Instruct model, after reconstructing the prompts for these 4 labels. Although there is still high amount of bias, we can filter them out and still have enough unbiased texts (more than 800 text passages per class).
To alleviate this issue, for the labels of man, woman, boy and girl, we reconstructed the prompts for text generation. The prompt used is the following:
Write two sentences with topic: <topic>, and keywords: <keyword1>, <keyword2>.
where <topic> was replaced with one of the following words: ‘factual’, ‘fiction’, ‘history’, ‘books’, ‘movies’, ‘philosophical’, and <keyword1> with one of the following words: ‘man’, ‘woman’, ‘boy’, ‘girl’. <keyword2> was treated differently depending on the label: For the term ‘man’, it was replaced with one of the following words:
actor, king, scientist, doctor, wizard, duke, lord, governor, prime minister, father, sorcerer, waiter, chess, director, producer, uncle, singer
For the term woman, it was replaced with one of the following words:
actress, queen, scientist, doctor, witch, duchess, lady, governor, prime minister, mother, sorcerer, waitress, chess, director, producer, aunt, singer
For the term boy, it was replaced with one of the following words:
kid, actor, prince, son, nephew, pupil, student, singer
And for the term girl, it was replaced with one of the following words:
kid, actress, princess, daughter, niece, pupil, student, singer
We then performed another round of bias detection using the Gemma model. While we found that a significant amount of bias still exists, we identified enough unbiased texts (according to the Gemma model) to include in the LUMA dataset. The number of biased and unbiased texts after re-generating the data for these 4 classes can be found in Figure 9.
To validate that the generated text passages could be correctly classified, we fed them back into the Gemma model using the following prompt:
In the text: "<text>", you need to understand what is the <masked_word> or find out what is the whole text about.
You need to classify the text into one of these classes: "<classes>"
[…]
% [TODO] Add audio generation pipeline
Appendix B Details about Baseline architectures and Datasets
Parameter Clean Dataset Reduced Diversity Increased Label Noise Increased Sample Noise
Compactness 0 20 0 0
Sample Noise False False False True
Label Switch Prob. 0 0 0.3 0
Noisy Data Ratio 0.0 0.0 0.0 1
Audio Noise SNR None None None 3-5
Image Noise None None None
Gaussian Noise: 4,
Shot Noise: 4,
Impulse Noise: 4,
Defocus Blur: 4,
Frosted Glass Blur: 4,
… (additional noises
summarized)
Text Noise None None None
KeyboardNoise: 1-5 char,
3-8 word,
BackTranslation,
Spelling: 0.4,
OCR: 0.5 word
Table 3: The parameters for generating the different versions of the dataset for experiments with the baseline models.
B.1 Dataset Generation
We compiled one clean version and three noisy versions of the dataset, each with a specific type of noise: reduced diversity, increased label noise, and increased sample noise. The parameters for compiling the different versions can be found in Table 3. These parameters were selected to introduce sufficient noise, ensuring observable changes in the uncertainty metrics. Please refer to our GitHub directory for the full configuration files.
B.2 Baseline models
In this section, we will describe the baseline models employed for evaluating the dataset. These baselines were selected to instantiate unimodal and multimodal architectures, which can be trained on the dataset and are not intended to serve as a comprehensive benchmark, nor did we endeavor to achieve the best possible performance.
For the image modality, we used a simple convolutional neural network, as depicted in Figure 10. For the audio modality, we extracted 128x128 mel-spectrograms from padded audio samples, and used a convolutional network for classification, as depicted in Figure 1. For the text modality, we extracted the BERT [5] embeddings for each token, and averaged them out, so that we have one embedding per text passage. Then, we passed the embedding through a simple feed-forward neural network (Figure 12) to get the predictions. As depicted in Figures 10, 11 and 12, each model includes two output heads: one for the prediction and the other for aleatoric uncertainty, following the methodology outlined by Valdenegro-Toro and Mori [34].
Figure 10: The classification network for the image modality.
Figure 11: The classification network for the audio modality.
Then, to combine the aforementioned unimodal networks into a multimodal architecture, we adopted the late fusion approach. In the Monte Carlo Dropout and Deep Ensemble methods, we obtained the multimodal prediction by averaging the logits from the final layers of the classifiers. For the RCML, we modified the output of the last layer in each network to produce evidence, as described in [37], and followed their methodology for combining the evidence.
In Figures 10, 11 and 12 the Dropout layer is kept turned on during the inference stage, for Monte-Carlo Dropout method, and turned off for Deep Ensemble and RCML methods. The dropout probability for all models is 0.3. The number of networks in deep ensembles is 10.
Figure 12: The classification network for the text modality.
We trained the networks for up to 300 epochs, with early stopping applied after 10 epochs of no improvement in validation loss. The learning rate was set to 0.001 and was reduced by a factor of 0.33 if there was no improvement in validation loss after 5 epochs.
The experiments were conducted on an internal cluster equipped with NVIDIA GeForce RTX 2080 GPUs, totaling approximately 22 GPU hours.
Appendix C Datasheet
In this section, we provide the datasheet [8] for LUMA dataset.
  1. 1.
    Motivation for Dataset Creation
    • Why was the dataset created? There is a lack of multimodal datasets that allow for controlled injection of different types of uncertainty for various data modalities. This dataset was created to address these needs and to provide a benchmark for learning from uncertain and multimodal data, offering a valuable resource to the research and R&D communities.
    • Has the dataset been used already? If so, where are the results so others can compare. We provide baseline results in this article, but the dataset has not been used yet in other research articles.
    • What (other) tasks could the dataset be used for? The dataset can be used to study the behavior of uncertainty quantification algorithms under different types of uncertainty in the data. The dataset can also be used for benchmarking multimodal deep learning models, robustness analysis, and to study modality imbalance [29] in multimodal learning.
    • Who funded dataset creation? This work is funded by Institut Archimedes doctoral program of Aix-Marseille University (AMU) in France.
  2. 2.
    Dataset Composition
    • What are the instances? In LUMA dataset, there are three modalities: image, audio, and text, belonging to 50 classes. Image modality contains images of different objects from CIFAR 10/100 classes [19] (e.g., horse, truck, man, bird, etc.). The audio modality contains audio tracks of people pronouncing these class labels. The text modality contains texts generated by Large Language Models (LLM) about the class labels.
    • How many instances are there in total (of each type, if appropriate)? In total, LUMA dataset has 135,127 audio samples, 62,875 text samples, 62,400 images from CIFAR-10/100, and around half a million generated images similar to CIFAR-100 data (belonging to the 50 classes). Nevertheless, LUMA dataset is a superset, and the open-source noise generation tool we provide takes a sample of 42 classes from each modality, with 500 training and 100 testing samples in each class. It also provides 3,859 Out-of-Distribution (OOD) samples.
    • What data does each instance consist of? Each instance consists of semantically related raw image, text and audio data.
    • Is there a label or target associated with each instance? Yes, each instance belongs one of 50 classes, which are selected from the CIFAR-10 and CIFAR-100 classes. The classes are: [‘man’, ‘boy’, ‘house’, ‘woman’, ‘girl’, ‘table’, ‘road’, ‘horse’, ‘dog’, ‘ship’, ‘bird’, ‘mountain’, ‘bed’, ‘train’, ‘bridge’, ‘fish’, ‘cloud’, ‘chair’, ‘cat’, ‘baby’, ‘castle’, ‘forest’, ‘television’, ‘bear’, ‘camel’, ‘sea’, ‘fox’, ‘plain’, ‘bus’, ’snake’, ‘lamp’, ‘clock’, ‘lion’, ‘tank’, ‘palm’, ‘rabbit’, ‘pine’, ‘cattle’, ‘oak’, ‘mouse’, ‘frog’, ‘ray’, ‘bicycle’, ‘truck’, ‘elephant’, ‘roses’, ‘wolf’, ‘telephone’, ‘bee’, ‘whale’].
    • Is any information missing from individual instances? There is no information missing in the LUMA dataset, but it is possible to reduce the information by the provided open-source noise generator tool to inject various types of uncertainty in the data.
    • Are relationships between individual instances made explicit? This is not applicable as we do not study the relationship between different data instances. Instead, we focus exclusively on the relationship between different modalities of the same instance, which are semantically related. For example, for the class cat, each instance contains an image of a cat, an audio pronouncing the word "cat", and a text passage about a cat.
    • Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set? The dataset is a superset containing all possible instances. The provided open-source tool shall be used to extract samples from the dataset with different amounts of uncertainties.
    • Are there recommended data splits (e.g., training, development/validation, testing)? If so, please provide a description of these splits, explaining the rationale behind them. Yes, the provided open-source tool extracts a recommended train and test sets with 600 and 100 samples respectively per class. For image data, we use the samples of the original train/test splits. For the other modalities, we split the data at random, trying to have a maximally diverse test set (see the main paper for the definition of diversity).
    • Are there any errors, sources of noise, or redundancies in the dataset? The text modality is generated through Gemma-7B Instruct [26] Large Language Model and can contain factual inaccuracies and biases. For the audio modality, we used automatic validation for 130,069 samples, as well as the help of 17 human annotators for manual validation of 8,372 samples.
    • Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)? The audio and text modalities are self-contained. For the image modality, due to the absence of an explicit license in the CIFAR dataset, we decided to provide the URL to the dataset and a code to download and align with the other modalities.
  3. 3.
    Collection Process
    • What mechanisms or procedures were used to collect the data? For the audio modality, three audio datasets (LibriSpeech [28], The Spoken Wikipedia [16] and Mozilla Common Voice [2]) were downloaded from their respective websites for later processing. For the image modality, CIFAR-10 and CIFAR-100 [19] datasets were downloaded from their website, and EDM [14] generated images were downloaded from https://huggingface.co/datasets/P2333/DM-Improves-AT. The text modality was generated using Gemma-7B Instruct [26] Large Language Model (LLM).
    • How was the data associated with each instance acquired? For the audio modality, each instance was obtained by extracting the pronunciation of specific words from audio/text parallel datasets using Algorithm 1. The audio samples were initially validated automatically using the Whisper speech-to-text model [31], followed by manual validation for the remaining samples. For the image modality, each instance was selected from a subset of images from the CIFAR-10 and CIFAR-100 datasets [19], as well as from images generated using the EDM generative model [14]. For the text modality, instances were generated using the Gemma-7B Instruct Large Language Model (LLM) [26] and automatically validated using the same language model.
    • If the dataset is a sample from a larger set, what was the sampling strategy The samples were selected to ensure an equal number of instances for each class. Additionally, based on user-specified parameters, the sampling method adjusts to include either highly diverse instances (uniform sampling) or less diverse ones, as detailed in Section 3.4.1.
    • Who was involved in the data collection process (e.g., students, crowdworkers, contractors) and how were they compensated (e.g., how much were crowdworkers paid)? Data collection was primarily carried out by the authors of this paper. For manual audio validation, our colleagues, including PhD and Master’s students as well as professors, volunteered to assist with the annotations. In total we had 16,744 annotations from 17 annotators.
    • Over what timeframe was the data collected? Does this timeframe match the creation timeframe of the data associated with the instances (e.g., recent crawl of old news articles)? If not, please describe the timeframe in which the data associated with the instances was created. The data was generated, collected, and validated between February 3, 2024, and May 5, 2024. The audio and image datasets were created much earlier. The Spoken Wikipedia dataset was released in 2017, the version of the Mozilla Common Voice dataset used was released on January 5, 2024, and the LibriSpeech dataset was released in 2015. The CIFAR datasets were released in 2009, and the EDM-generated images were released in February 2023.
  4. 4.
    Data Preprocessing
    • Was any preprocessing/cleaning/labeling of the data done? Yes, data preprocessing, cleaning and validation was performed for audio and text modalities, as described in Section 3.
    • Was the "raw" data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the "raw" data. The "raw" data is not saved, but audio and image modalities are available online at the links provided in the main paper.
    • Is the software used to preprocess/clean/label the instances available? If so, please provide a link or other access point. The source code for data acquisition, pre-processing, cleaning and validation is not available. We do provide the source code for compiling the final versions of the LUMA dataset.
    • Does this dataset collection/processing procedure achieve the motivation for creating the dataset stated in the first section of this datasheet? If not, what are the limitations? Yes, the dataset allows to insert different types of noises to each of the modalities in a controlled manner.
  5. 5.
    Dataset Distribution
    • How will the dataset be distributed? (e.g., tarball on website, API, GitHub; does the data have a DOI and is it archived redundantly? The dataset is released in HuggingFace platform with DOI: 10.57967/hf/2502. The source code for LUMA uncertainty generation tool is open-sourced and available at https://github.com/bezirganyan/LUMA.
    • When will the dataset be released/first distributed? What license (if any) is it distributed under? The dataset is already distributed under the CC BY-SA 4.0 license.
    • Are there any copyrights on the data? The source audio datasets have CC BY-SA 4.0, CC BY 4.0 and CC0 licenses, and our use of the datasets corresponds to the terms in the licenses. The EDM Generated images have Apache-2.0 license. We could not find liceneses for the CIFAR datasets, hence, we merely provide links and source code to download these datasets from their official website.
    • Are there any fees or access/export restrictions? No
  6. 6.
    Dataset Maintenance
    • Who is supporting/hosting/maintaining the dataset? The dataset is hosted in the HuggingFace platform and will be maintained by the authors of the paper.
    • Will the dataset be updated? If so, how often and by whom? The dataset may be updated with more modalities by the authors of the paper. We will also accept contributions from the community.
    • How will updates be communicated? (e.g., mailing list, GitHub) The updated will be made explicit in the Dataset Card in the HuggingFace platform.
    • If the dataset becomes obsolete how will this be communicated? Any changes to the dataset or its relevance will be communicated in its HuggingFace page.
    • Is there a repository to link to any/all papers/systems that use this dataset? No
    • If others want to extend/augment/build on this dataset, is there a mechanism for them to do so? If so, is there a process for tracking/assessing the quality of those contributions. What is the process for communicating/distributing these contributions to users? We welcome contributions from the community. The contributors are welcome to open issues or pull requests in our GitHub and HuggingFace pages. If the contributions align with the purpose of the dataset, the authors will merge them into the respective repositories.
  7. 7.
    Legal and Ethical Considerations
    • Were any ethical review processes conducted (e.g., by an institutional review board)? No
    • Does the dataset contain data that might be considered confidential? No
    • Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? Yes. The text modality was generated using the Gemma-7B instruct large language model, and hence, inherited the biases and stereotypes form the model. While we performed an automatic bias detection procedure as described in Section 3.3, there may still be samples that were not detected by our bias detection algorithms.
    • Does the dataset relate to people? No